News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 30 August 2002 10.00am EDT

During the last week the observing schedule was interrupted following a high radiation event at 9:11pm EDT on Aug 23 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and radiation levels dropped sufficiently to allow the loads to be restarted at 8:26am on Aug 25 in time to perform an ACA dark current calibration. A total of 126.9Ks of observing time was impacted by the event and the affected observations will be rescheduled in a later load.

Preparations are complete for a 39.5 minute duration lunar eclipse that will occur at 11:55am on Sep 7. Software patches to provide the appropriate power system configuration and provide for contingencies will be uplinked on Sept 6.

The observing schedule for next week is shown below and includes a roll constrainted observation of Cen-A.

------------------------------------------
Radiation Belts                      Sep 2
Cen-A                   ACIS-S
Aql X-1                 ACIS-S       Sep 3
RXJ1856.5-3754          HRC-I
2S 1845-024             ACIS-S/HETG  Sep 4
Zeta Oph                ACIS-S/HETG
Radiation Belts                      Sep 5
Zeta Oph                ACIS-S/HETG
2S 1845-024             ACIS-S/HETG  Sep 6
ONC Flanking Field S    ACIS-I
A68                     ACIS-I
MACSJ2311.5+0338        ACIS-I       Sep 7
PSS 0008+1923           ACIS-S
Radiation Belts
PSS0808+5215            ACIS-S       Sep 8
PSS 1507+5603           ACIS-S
AX J1531.8+2414         ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports